# Position Paper: Rethinking Empirical Research in Machine Learning: Addressing Epistemic and Methodological Challenges of Experimentation

### v1

@@ 3 Improving the Status Quo: More Richness in Empirical Methodological Research
Moreover, without more education, none of this will be possible. Given the different perspectives – formal science, engineering, statistical – from which ML can be viewed, it is very difficult to include each in the appropriate depth in a single study program. While a recent survey of 101 undergraduate data science programs in the U.S. showed that all included an introductory course in statistics (Bile Hassan et al., 2021), statistics has only recently (2023) been included as a core topic in the curriculum recommendations for CS (Joint Task Force on Computing Curricula, 2023). It is also questionable if introductory courses are sufficient to avoid crucial gaps that can lead to the adoption of questionable research practices (cf. Gigerenzer, 2018). Furthermore, nearly no study program contains a dedicated course on Design and Analysis of (Computer) Experiments (Santner et al., 2003; Box et al., 2005; Dean et al., 2017), which we deem especially relevant for our context here. In general, we agree with De Veaux et al. (2017, p. 16-17) that many “courses traditionally found in computer science, statistics, and mathematics offerings should be redesigned for the data science [or ML] major in the interest of efficiency and the potential synergy that integrated courses would offer.”
@@
Advice for junior researchers. (1) Read the positive examples of insight-oriented exploratory research in ML (listed above), about the design of experiments, the critical discussion on statistical testing, and the basics of philosophy of science. (2) Educate yourself in Open Science practices (e.g., see The Turing Way Community, 2023). (3) Engage with researchers from other disciplines as data (the one which ML models are trained on) can only really be understood if one understands how it was generated. (4) Consider making empirical research in ML a (partial) research focus.
Advice for senior researchers. (1) Allow your junior researchers to write (great) papers on empirical aspects of ML, even if those may be relatively difficult to publish in major venues for now. Our personal experience is that these papers can still be highly cited and become very influential. (2) Learn from other fields; what we are experiencing in terms of non-replicable research is not a new phenomenon. (3) Please do not perceive this paper as an ‘attack’ on ML but rather as an honest attempt to improve it and, more importantly, to improve its impact.
Advice for venue organizers and editors (see also Nakkiran & Belkin, 2022). (1) Encourage all forms of proper empirical ML to be submitted (in particular, this includes insight-oriented exploratory research), e.g., by creating special tracks or adding keywords but also by allowing such work on ‘main tracks’. The idea is to create special measures for the topic to increase awareness but not to ‘isolate’ or ‘ban’ all such papers to special (workshop) tracks with (potentially) ‘less’ value. (2) Consider giving out awards for positive examples of these types of research. (3) Consider establishing positions like reproducibility and replicability editors for venues and journals. (4) Give concrete advice for best practices, so authors and reviewers have clear guidelines to follow. Note that this should not be confined to asking ‘Were the empirical results subjected to statistical tests?’ (without further information); this is close to the opposite of what we think is needed.999 One anonymous reviewer also suggested that venues start collecting meta data on reasons for rejection. Such data could serve as a basis to evaluate if certain types of ML research face a systematic bias.
@@ 4 Beyond the Status Quo: Rethinking Empirical ML as a Maturing Science
Based on that notion, Fife & Rodgers (2022) argue that “psychology may not be mature enough to justify confirmatory research” (p. 453) and that “[t]he maturity of any science puts a cap on the exploratory/confirmatory continuum” (p. 462). Given the similarities between research in psychology and ML as described by Hullman et al. (2022), we think similar holds for ML and we suggest ML should be considered as a maturing (empirical) science as well. [^10] There are also differences between ML and psychology considerably simplifying our lives: we usually do not experiment on humans but algorithms on computers and have more control over experiments, larger sample sizes, and lower experimental costs. Hullman et al. (2022, p. 355) “identify common themes in reform discussions, like overreliance on asymptotic theory and non-credible beliefs about real-world data-generating processes.” That said, confirmatory research in ML as advocated in Section 3 is still very different from strict confirmatory research in other disciplines. Rather it can be seen as “rough” confirmatory research (Fife & Rodgers, 2022; Tukey, 1973) that follows the same principles, but – as outlined – it is unclear how results can be generalized (e.g., using statistical tests) – a cornerstone of strict confirmatory research. But this should not be taken as a caveat. Rough confirmatory research allows for flexibility that strict confirmatory research does not (Fife & Rodgers, 2022).
@@
Statistical significance tests: Words of caution, revisited! The problem of empirical research as a continuum is more involved epistemologically and cannot be discussed in full detail here. An important aspect that needs to be discussed is its relation to the misguided use of statistical testing. This point has been made by Drummond (2006) before and in more detail. We revisit it here, enriching it with more recent literature on the issue. In particular, routinely adding statistical machinery to an (already underspecified and/or biased) experimental design to test for statistically significant differences in performance – as is frequently done and/or explicitly asked for (e.g., Henderson et al., 2018; Marie et al., 2021) – does not improve the epistemic relevance of the results by much nor does it add much additional insight over other data aggregations. In fact, “[s]tatistical significance was never meant to imply scientific importance”, and you should not “conclude anything about scientific or practical importance based on statistical significance (or lack thereof)” (Wasserstein et al., 2019, pp. 2, 1). On the contrary, the misguided beliefs in and use of statistical rituals (Gigerenzer, 2018) is largely responsible for the replication crisis in other empirical disciplines.
The reasons are complex. First of all, the modern theory of statistical hypothesis testing (SHT) is a conflation of two historically distinct types of testing theory. Important epistemological questions about when statistical tests are appropriate are obscured by this mixed theory (e.g., Schneider, 2015; Gigerenzer & Marewski, 2015; Rubin, 2020). More importantly, specifically for experiments in ML, both theories are developed for experimental designs based on samples randomly drawn from a population of interest (Schneider, 2015). In general, the assumptions underlying the theory of statistical testing as an inferential tool are usually not met in many applications (Greenland, 2023). In fact, the editors of the The American Statistician special issue Statistical Inference in the 21st Century: A World Beyond $p<0.05$ went so far as to conclude, “based on [their] review of the articles in this special issue and the broader literature, that it is time to stop using the term ‘statistically significant’ entirely” (Wasserstein et al., 2019, p. 2).
@@
Most current empirical ML research should rather be viewed as exploratory.    As outlined, confirmatory research aims to test preexisting hypotheses, while exploratory research involves gathering information, identifying patterns, and formulating specific hypotheses for further investigation. We think, currently, most of the empirical research in ML is conducted as part of a paper introducing a new method and is fashioned as confirmatory research even though it is exploratory in nature. In our view, this is reflected especially in the routine use of statistical tests to aggregate benchmark results: the exploratory phase of method development (e.g., trying out different method variants) largely invalidates post-hoc statistical tests. As Strobl & Leisch (2024, p. 2) put it, “In methodological research, comparison studies are often published either with the explicit or implicit aim to promote a new method by means of showing that it outperforms existing methods.” In other words, the conducted experiments are set up to confirm the (implicit) hypothesis that the proposed method constitutes an improvement. Systemic pressures and conventions, as well as ML’s strong roots in formal sciences and focus on improving applications, encourage this mindset and the practice of invoking confirmatory arguments. This is expressed in statements such as that “[i]t is well-known that reviewers ask for application improvements” and “for ‘theoretical justification’ for purely experimental papers, even when the experiments alone constitute a valid scientific contribution” (Nakkiran & Belkin, 2022, pp. 2-3). The problem with not emphasizing the exploratory nature is that “exploratory findings have a slippery way of ‘transforming’ into planned findings as the research process progresses” (Calin-Jageman & Cumming, 2019, p. 275) and “[a]t the bottom of that slippery slope one often finds results that don’t reproduce” (Wasserstein et al., 2019, p. 3). Shifting the focus to an exploratory notion of method development is an opportunity to fully allow “to understand under which circumstances the algorithm produces misleading results” (Rendsburg et al., 2020, p. 9) and to “learn about [its] strengths and weaknesses” (Sculley et al., 2018, p. 2) and clearly report them.

### v2

@@ 3 Improving the Status Quo: More Richness in Empirical Methodological Research
Moreover, without more education, none of this will be possible. Given the different perspectives – formal science, engineering, statistical – from which ML can be viewed, it is very difficult to include each in the appropriate depth in a single study program. While a recent survey of 101 undergraduate data science programs in the U.S. showed that all included an introductory course in statistics (Bile Hassan et al., 2021), statistics has only recently (2023) been included as a core topic in the curriculum recommendations for CS (Joint Task Force on Computing Curricula, 2023). It is also questionable if introductory courses are sufficient to avoid crucial gaps that can lead to the adoption of questionable research practices (cf. Gigerenzer, 2018). Furthermore, nearly no study program contains a dedicated course on design and analysis of (computer) experiments (Santner et al., 2003; Box et al., 2005; Dean et al., 2017), which we deem especially relevant for our context here. In general, we agree with De Veaux et al. (2017, pp. 16–17) that many “courses traditionally found in computer science, statistics, and mathematics offerings should be redesigned for the data science [or ML] major in the interest of efficiency and the potential synergy that integrated courses would offer.”
@@
Advice for junior researchers. (1) Read the positive examples of insight-oriented exploratory research in ML (listed above), about the design of experiments, the critical discussion on statistical testing, and the basics of philosophy of science. (2) Educate yourself in Open Science practices (e.g., see The Turing Way Community, 2023). (3) Engage with researchers from other disciplines as data (the one on which ML models are trained) can only really be understood if one understands how it was generated. (4) Consider making empirical research in ML a (partial) research focus.
Advice for senior researchers. (1) Allow your junior researchers to write (great) papers on empirical aspects of ML, even if those may be relatively difficult to publish in major venues for now. Our personal experience is that these papers can still be highly cited and become very influential. (2) Learn from other fields; what we are experiencing in terms of non-replicable research is not a new phenomenon. (3) Please do not perceive this paper as an attack on ML but rather as an honest attempt to improve it and, more importantly, to improve its impact.
Advice for venue organizers and editors (see also Nakkiran & Belkin, 2022). (1) Encourage all forms of proper empirical ML to be submitted (in particular, this includes insight-oriented exploratory research), e.g., by creating special tracks or adding keywords but also by allowing such work on main tracks. The idea is to create special measures for the topic to increase awareness but not to isolate or ban all such papers to special (workshop) tracks with (potentially) lower perceived value. (2) Consider giving out awards for positive examples of these types of research. (3) Consider establishing positions like reproducibility and replicability editors for venues and journals. (4) Give concrete advice for best practices, so authors and reviewers have clear guidelines to follow. Note that this should not be confined to asking “Were the empirical results subjected to statistical tests?” (without further information); this is close to the opposite of what we think is needed. [^9] One anonymous reviewer also suggested that venues start collecting metadata on reasons for rejection. Such data could serve as a basis to evaluate if certain types of ML research face a systematic bias.
@@ 4 Beyond the Status Quo: Rethinking Empirical ML as a Maturing Science
Based on that notion, Fife & Rodgers (2022) argue that “psychology may not be mature enough to justify confirmatory research” (p. 453) and that “[t]he maturity of any science puts a cap on the exploratory/confirmatory continuum” (p. 462). Given the similarities between research in psychology and ML as described by Hullman et al. (2022), we think similar holds for ML and we suggest ML should be considered as a maturing (empirical) science as well. [^10] There are also differences between ML and psychology considerably simplifying our lives: we usually do not experiment on humans but algorithms on computers and have more control over experiments, larger sample sizes, and lower experimental costs. Hullman et al. (2022, p. 355) “identify common themes in reform discussions, like overreliance on asymptotic theory and non-credible beliefs about real-world data-generating processes.” That said, confirmatory research in ML as advocated in Section 3 is still very different from strict confirmatory research in other disciplines. Rather it can be seen as rough confirmatory research (Fife & Rodgers, 2022; Tukey, 1973) that follows the same principles, but – as outlined – it is unclear how results can be generalized (e.g., using statistical tests) – a cornerstone of strict confirmatory research. But this should not be taken as a caveat. Rough confirmatory research allows for flexibility that strict confirmatory research does not (Fife & Rodgers, 2022).
@@
Statistical significance tests: Words of caution, revisited! The problem of empirical research as a continuum is more involved epistemologically and cannot be discussed in full detail here. An important aspect that needs to be discussed is its relation to the misguided use of statistical testing. This point has been made by Drummond (2006) before and in more detail. We revisit it here, enriching it with more recent literature on the issue. In particular, routinely adding statistical machinery to an (already underspecified and/or biased) experimental design to test for statistically significant differences in performance – as is frequently done and/or explicitly asked for (e.g., Henderson et al., 2018; Marie et al., 2021) – does not improve the epistemic relevance of the results by much nor does it add much additional insight over other data aggregations. In fact, “[s]tatistical significance was never meant to imply scientific importance,” and you should not “conclude anything about scientific or practical importance based on statistical significance (or lack thereof)” (Wasserstein et al., 2019, pp. 2, 1). On the contrary, the misguided beliefs in and use of statistical rituals (Gigerenzer, 2018) is largely responsible for the replication crisis in other empirical disciplines.
The reasons are complex. First of all, the modern theory of statistical hypothesis testing (SHT) is a conflation of two historically distinct types of testing theory. Important epistemological questions about when statistical tests are appropriate are obscured by this mixed theory (e.g., Schneider, 2015; Gigerenzer & Marewski, 2015; Rubin, 2020). More importantly, specifically for experiments in ML, both theories are developed for experimental designs based on samples randomly drawn from a population of interest (Schneider, 2015). In general, the assumptions underlying the theory of statistical testing as an inferential tool are usually not met in many applications (Greenland, 2023). In fact, the editors of the The American Statistician special issue “Statistical Inference in the 21st Century: A World Beyond $p<0.05$” went so far as to conclude, “based on [their] review of the articles in this special issue and the broader literature, that it is time to stop using the term ‘statistically significant’ entirely” (Wasserstein et al., 2019, p. 2).
@@
Most current empirical ML research should rather be viewed as exploratory.    As outlined, confirmatory research aims to test preexisting hypotheses, while exploratory research involves gathering information, identifying patterns, and formulating specific hypotheses for further investigation. We think, currently, most of the empirical research in ML is conducted as part of a paper introducing a new method and is fashioned as confirmatory research even though it is exploratory in nature. In our view, this is reflected especially in the routine use of statistical tests to aggregate benchmark results: the exploratory phase of method development (e.g., trying out different method variants) largely invalidates post hoc statistical tests. As Strobl & Leisch (2024, p. 2) put it: “In methodological research, comparison studies are often published either with the explicit or implicit aim to promote a new method by means of showing that it outperforms existing methods.” In other words, the conducted experiments are set up to confirm the (implicit) hypothesis that the proposed method constitutes an improvement. Systemic pressures and conventions, as well as ML’s strong roots in formal sciences and focus on improving applications, encourage this mindset and the practice of invoking confirmatory arguments. This is expressed in statements such as that “[i]t is well-known that reviewers ask for application improvements” and “for ‘theoretical justification’ for purely experimental papers, even when the experiments alone constitute a valid scientific contribution” (Nakkiran & Belkin, 2022, pp. 2–3). The problem with not emphasizing the exploratory nature is that “exploratory findings have a slippery way of ‘transforming’ into planned findings as the research process progresses” (Calin-Jageman & Cumming, 2019, p. 275) and “[a]t the bottom of that slippery slope one often finds results that don’t reproduce” (Wasserstein et al., 2019, p. 3). Shifting the focus to an exploratory notion of method development is an opportunity to fully allow “to understand under which circumstances the algorithm produces misleading results” (Rendsburg et al., 2020, p. 9) and to “learn about [its] strengths and weaknesses” (Sculley et al., 2018, p. 2) and clearly report them.